Suppose there exist 2 countries, Home and Foreign; 2 goods, X and Y; and 2 factors of production, labor (L) and capital (K). Each country can produce both goods. X is labor-intensive and Y is capital-intensive. Home is labor-abundant and Foreign is capital-abundant. Assume that the standard assumptions of the Heckscher-Ohlin model hold. When answering the following question, please support each of your arguments with detailed analysis and draw the relevant diagrams to support your answer. Consider a move from autarky to free trade between the two countries in both goods. Identify which group of people (laborers or owners of capital) gain and which lose in Home from the move from autarky to free trade. Discuss clearly how the assumptions of the model lead to this result.
Answer #1

In a world of two commodities(X and Y) and two factors (labor and Capital ),we say that commodity Y is capital intensive if the capital labor ratio (K/L)used in the production of Y is greater than K/L used in the production of X.

Assumptions of Heckscher-Ohlin theory:

1.Two nations,Two commodities,and Two Factors.

2.Both nations use same technology

3.Commodity X is labor intensive and commodity Y is capital intensive.

4.Constant returns to scale in the production of both commodities in both nations.

5.Incomplete specialisation in production in both nations.

6.Tastes are equal in both nations.

7.Perfect competition in both commodities and factor markets in both nations.

8.There is perfect factor mobility within each nation but no international factor mobility.

9.There are no transportation costs,tariffs,or other obstructions to the free flow of international trade.

10.All resources are fully employed in both nations.

11.International trade between two nations is balanced.

Nation 2 in 8=4 Nation 1 • 2Y Ain V=1 1Y 27 2x - 4 in x = 1/2 2X 1X LLLL 0 2 4 6 8 10 12 0 1 2 4 6

Factor intensities for commodity X and Y in Nation 1 and Nation 2

In nation 1,the capital labor ratio(K/L)equals 1 for commodity Y and K/L=1/4 for commodity X.These are given by the slope of the ray from the origin for each commodity in Nation 1.Thus commodity Y is K intensive commodity in Nation 1.In Nation 2,K/L=4 for Y and K/L=1 for X.Thus commodity Y is K intensive and commodity X is L intensive in both countries.Nation 2 uses higher K/L than nation 1 in the production of both commodities,because the relative price of capital is lower in Nation 2.

140 Nation 2 120 100 80 Nation 1 0 20 40 60 80 100 120 140